OpenStudy (anonymous):

According to the tan graph , 1/0 = infinity , would that mean 0 * infinity = 1 ??

ganeshie8 (ganeshie8):

good q :) honestly i dont have perfect explanation, but let me ask you a q... Is the slope of a vertical line infinity or undefined ?

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

no i don't think so...

OpenStudy (anonymous):

Can u say whats the tan 90 ? We cant say , its vertical and in another language a line make 90 dragee angle with X s is a Vertical line :) Just this :) Got it ?

OpenStudy (anonymous):

"infinity" is nothing! so i think we can't say that...

OpenStudy (anonymous):

But I think Infinity is a thing and we can say that but we cant find it :)

OpenStudy (anonymous):

if infinity is a thing so 0 * infinity is 0 not 1 :)

OpenStudy (anonymous):

so it seems not correct

OpenStudy (anonymous):

@ganeshie8 , isn't the tan graph ever closing in on the y-axis but never quite reaching ? Wouldn't that mean in fact that it is infinity ?

OpenStudy (anonymous):

Yes ! Would mean :)

OpenStudy (anonymous):

So... , wouldn't that make 0* infinity = 1 , arithmetically

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

@E.ali SO you're saying that if we keep adding zero (nothing) , we eventually reach 1?

OpenStudy (anonymous):

No ! I just said we can say Infinity is a think that we didnt know that :)

ganeshie8 (ganeshie8):

if u keep adding 0, you may run into debt becoz of ever increasing inflation :)

ganeshie8 (ganeshie8):

actually i like how u put it above :- ... isn't the tan graph ever closing in on the y-axis but never quite reaching ? Wouldn't that mean in fact that it is infinity ?

ganeshie8 (ganeshie8):

we observe that, tan(90) is approaching y-axis, you mean it is ever increasing/ ever decreasing eh ? (+infinity or - infinity)

OpenStudy (anonymous):

guys , gusy! remember! "infinity" is not a thing that follows these rules and laws!

ganeshie8 (ganeshie8):

yes infinity is not an useful thing, or not even a thing in algebra but its a very useful concept (ok i agree its not a thing :)) in calculus

ganeshie8 (ganeshie8):

in algebra+trig, \(\large \frac{1}{0} = UNDEFINED\)

OpenStudy (anonymous):

In calculus !

ganeshie8 (ganeshie8):

in calculus, \(\large as\ x-> 0, \frac{1}{x} -> \infty\) as something approaches something, something else happens... thats how it works wid infty in calculus

ganeshie8 (ganeshie8):

ur actual q, 1/0 = infinity , would that mean 0 * infinity = 1 the moment u see 1/0, just shout UNDEFINED and stop. dont mess wid it.

OpenStudy (anonymous):

\[ x\to 0\implies \left|\frac 1x \right|\to \infty \]You can't leave out the absolute value bars.

ganeshie8 (ganeshie8):

yup they're mandatory

OpenStudy (anonymous):

HAHAHAA , this was a lot of fun , guys, I just wanted to see how long it would take to reach the UNDEFINED part :D , @ganeshie8 , u da MAN , THANKS a lot @wio , @E.ali , @PFEH.1999

ganeshie8 (ganeshie8):

lol i sensed it, u asked d q purposefully to stir some discussion nice :)

OpenStudy (akashdeepdeb):

The problem with the question is that 1/0 is not actually infiniti It is approaching infiniti And on top of that Infiniti is JUST AN IDEA and not a specific number or a quantity. 1/0 = IDEA :D LOL :')